wiki:UC20210105

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 01/05/2021

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Denise Stephens (BYU), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Kal Kadlec (UW), Chip Kobulnicky (UWy), Bill Ketzeback (APO), Russet McMillan (APO), Jeff Bary (NAPG), Sarah Tuttle (UW)


User feedback and comments from institutional representatives

  • FGCU - nothing to report, other than observing has been going well (modulo the weather)
  • Seattle - nothing to report
  • Washington - nothing to report
  • NAPG - nothing to report
  • JHU - nothing to report
  • UVa - nothing to report
  • Georgia State - nothing to report
  • Oklahoma - a user had a few issues lately (related to the sticky ARCTIC shutter and guider problems) that were not mentioned in the Instrument News: Current Status page on the web site. We will make every attempt to keep the information on that page as up-to-date as possible.
  • BYU - nothing to report
  • NMSU - nothing to report
  • Wyoming - nothing to report
  • CU - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 12/05/2020 - 1/04/2021

0) Overview

APO has been operating at the 25% staffing limit since reopening on 12/2/2020. The State of New Mexico updated the tiered positions for every county on 12/16 and 12/30; Otero County (where APO is located) was in the red tier for both updates. The numbers for the 12/30 update were much better though: 6.44% positivity rate and 23 cases per 100k people. At this rate of improvement, it is very possible that on the 1/13/21 update, Otero County will move to the yellow tier (although that would not change staffing levels for APO, since those don’t change until reaching the green tier).  For comparison, the numbers for Dona Ana County (where NMSU is located) from the last state update were 12.82% positivity rate and 43.7 cases per 100k.

The data and numbers are available here: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/ 

We are still restricting all visitors to the site. Following the State of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

Weather for the month of December has been colder than average; precipitation amounts are below average also.

As some of you may know, 3.5m Observing Specialist Ted Rudyk is leaving APO for Gemini North. With Ted leaving, we will be short handed on the 3.5m night staff until his replacement is hired and trained.  This staffing shortage may cause a slow-down in response time; we appreciate everyone's understanding during this time.

1) Telescope

The telescope has been performing as expected.

2) Instruments

DIS: The scattered light in both the red and the blue cameras is stable and considered good to excellent.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light ratio (IOL) for red and blue continues to improve slowly with cooling weather.

NIC-FPS: powered on, cold and operational.

ARCTIC: powered on, cold, and operational. The issue with dome flats is being debugged and solutions are being worked on.  The design for the “fence” baffle on the mirror cover petal is being worked on; initial design ideas had to be modified because of interference issues with the telescope structure. The new redesigned rotation mechanism for the diffuser has arrived on site and installation will begin when staffing levels permit.  As a reminder, the ARCTIC shutter has issues when the temp approaches -10 degrees C. 

3) ARCSAT 

ARCSAT is operational and has been used for observations this past month. Driver and software issues associated with changing temperature as well as upgraded software appear to be resolved.  The issue between the two modes of operations for pointing is still under investigation. The Shelyak eShel preliminary commissioning took place at the beginning of November.  The data from the commissioning work are still being analyzed and discussed. Further engineering is scheduled for Q1.


Additional telescope and instrument discussion

Given that we will be short-handed for nighttime observing staff for the near future, users are asked to respond to the reminder emails promptly so the site staff has plenty of lead time to set up the instrument(s) for a given night of observing. The job ad for a new observing specialist will be posted this week. The secondary mirror recently had some trouble moving from the focus position for one instrument to the focus position for another instrument (this happened on a particularly cold night). We think the problem has been fixed but it's something to watch out for. Some more work on testing the guiding for ARCSAT has been completed and the data are being evaluated.


APO and the COVID-19 virus

There is not much new to say other than what was covered by Jamey's report. The travel restrictions still impact class visits and student training opportunities, but we are now equipped to conduct those virtually. There are two training runs scheduled for Q1: CU and UWy (a class and a group of ~ 5 graduate students, respectively). There were no requests for a pick-up training run (for non-students) in Q1.


Q1 3.5m scheduling

There is some open time available in Q1: Jan 19 is the earliest slot, there are some dates in bright time, and the earliest A half is in mid-Feb. Interested users should also check the availability of DD01 time in Q1. Some of these slots are already claimed but not all; it doesn't hurt to ask. Many of the slots are short or not adjacent to twilight (those are particularly well-suited for echelle observations since the instrument has internal calibrations and doesn't require twilight time to do cals).


Q1 ARCSAT scheduling

The Q1 schedule is relatively full. There are 3 new ARCSAT observers this quarter, which is exciting. This week is listed as open but is planned for engineering; the end of March is also available. Additionally, there are some slots in February that are available (portions of several weeks), but they have to be with FlareCam.


KOSMOS Update

The team is regrouping after the holiday break; next tasks are to install the camera in the dewar and then conduct in-lab calibration and testing. In addition to characterizing the optical performance of the instrument in the lab, some code wrapping needs to be completed so that the instrument interfaces nicely with TUI.


ARCTIC Diffuser Rotation

Bill removed the diffuser from ARCTIC to start work on upgrading it for the rotation mechanism. The parts that need to be machined are in the shop and are being worked on today. ARCTIC is heavily scheduled throughout January so there are only a few short windows during which work on the electronics for the robotics can be completed. Bill will provide us with another update at our next meeting.


All Other Business

John Bally inquired about the status of the DAWG final report. Nancy reviewed a few of the recommendations, which fell into the general categories of data storage/management and instrument-related software. There was some discussion of software tools that have been developed for reduction of spectroscopic data from other observatories (e.g. PypeIt, https://arxiv.org/abs/2005.06505) and whether they could be adapted for ARCES purposes (and by whom). Nancy asked the UC reps to share the DAWG final report (which she will distribute) with their users and solicit feedback on which items are highest priority so that she can allocate resources most effectively.


ACTION ITEMS

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. A vendor has been decided on, and she will now request official quotes rather than rough order of magnitude cost estimates so that we can move forward with the purchase.
  • Develop a survey to get feedback on the pickup training session - Nancy. Status: OPEN. This is nearly done. Once it is completed, we will send it to those who were trained virtually in Q3 and Q4, as well as to the User Committee reps for distribution.

New action items from this meeting:

  • Send DAWG report to all UC reps - Nancy (note: it is also posted on the 3.5m users' wiki!). Status: CLOSED.
  • Distribute DAWG report to all 3.5m users - UC reps. Please solicit their feedback and be prepared to discuss this at our next meeting.

Next meeting

The next meeting will be on Tuesday February 2 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.